20men can do 1/6 work in 5 days. After working for how many days should 5 men leave the job so that 15 men will complete whole work in 35 days.
1
Expert's answer
2012-09-28T07:51:23-0400
There needed

N = 35*15 = 525 human days

to complete the work. So, 5 men will do that work in

M = 525/5 = 105 days.